# Customizing Language Models with Instance-wise LoRA for Sequential Recommendation

### v1

@@ Abstract
Sequential recommendation systems predict a user’s next item of interest by analyzing past interactions, aligning recommendations with individual preferences. Leveraging the strengths of Large Language Models (LLMs) in knowledge comprehension and reasoning, recent approaches have applied LLMs to sequential recommendation through language generation paradigms. These methods convert user behavior sequences into prompts for LLM fine-tuning, utilizing Low-Rank Adaptation (LoRA) modules to refine recommendations. However, the uniform application of LoRA across diverse user behaviors sometimes fails to capture individual variability, leading to suboptimal performance and negative transfer between disparate sequences. To address these challenges, we propose Instance-wise LoRA (iLoRA), integrating LoRA with the Mixture of Experts (MoE) framework. iLoRA creates a diverse array of experts, each capturing specific aspects of user preferences, and introduces a sequence representation guided gate function. This gate function processes historical interaction sequences to generate enriched representations, guiding the gating network to output customized expert participation weights. This tailored approach mitigates negative transfer and dynamically adjusts to diverse behavior patterns. Extensive experiments on three benchmark datasets demonstrate the effectiveness of iLoRA, highlighting its superior performance compared to existing methods in capturing user-specific preferences and improving recommendation accuracy.
@@ 1 Introduction
Sequential recommendation [1, 2, 3, 4, 5] suggests a user’s next item of interest by analyzing his/her past interactions, tailoring recommendations to individual preferences. As Large Language Models (LLMs) [6, 7, 8, 9, 10, 11, 12] exhibit impressive proficiency in global knowledge comprehension and reasoning, their potential for application in sequential recommendation is garnering increasing interest [13, 14, 15, 16, 17, 18, 19, 20]. Recent efforts [21, 22] approach the sequential recommendation task under a language generation paradigm, wherein user behavior sequences are converted into input prompts by either purely textual prompting (ID numbers or descriptions) [14, 16, 23, 24, 17] or hybrid prompting with additional behavioral tokens [21, 22, 25, 17], achieving remarkable success.
@@
Upon scrutinizing prior studies on LLM-based sequential recommenders, we can summarize a common fine-tuning pipeline comprising three sequential components: (1) convert a sequence of historical behaviors into a prompt; (2) pair these prompts with the subsequent items of interest, to create instruction-tuning datasets; (3) incorporate a trainable Low-Rank Adaptioin (LoRA) module [16, 25, 22, 21] into LLMs and fine-tune it on such prompts. Existing studies [26, 14, 24, 17, 21, 22, 25] primarily focus on refining high-quality prompts to more effectively incorporate recommendation information, while leaving the improvements to LoRA largely under-explored. Instead, they employ a single, standard LoRA module for fine-tuning, which freezes the LLM weights and updates the model through two additional trainable low-rank matrices.
@@
However, as user behaviors often exhibit substantial individual variability (e.g., distinct interests, behavior patterns, feedback mechanisms), we argue that employing a single, standard LoRA module across such diverse behaviors may not effectively capture these variabilities. Specifically, the inherent variability in user behaviors naturally inspires us to view item sequences with different behavior variables as different tasks. Simply applying a singular LoRA module leaves this multitask nature of sequential recommendation untouched, thus potentially overestimating the relationships between sequences. It easily causes the negative transfer between significantly discrepant sequences [27, 28]. Take LLaRA [21] as an example, which fine-tunes a LoRA module on top of LLaMA-2 [9] across all sequences. Following prior studies [28, 27], we analyze the gradients of LLaRA associated with disparate sequences and reveal significant misalignment, as shown in Figure 1. Specifically, we observe strong clustering by membership closeness in the collaborative space, along the diagonal of the gradient similarity matrix. Clearly, such gradients are misaligned, which can result in suboptimal performance. To mitigate this issue, one straightforward solution is to deploy multiple LoRA modules, each fine-tuned for a specific sequence, enabling each module to act as a lightweight expert tailored to its respective sequence. However, it is impractical in terms of resources and time complexity, as the number of sequences often scales to millions.
@@
Figure 1: Gradient similarity of LoRA modules across training steps, with sequence data partitioned into 8 clusters. Euclidean distance is employed to evaluate the proximity between clusters, whereas gradient similarity is measured to assess geometric loss. Darker cells indicate higher gradient similarity. A case study shows user sequences: those with 0.86 cosine similarity share a strong interest in thriller movies, while those with -0.75 cosine similarity lack noticeable preference similarities. Similar items are connected with dashed lines, while identical items are connected with solid lines.
@@ 2 Preliminary
Fine-tuning with Low-rank Adaption (LoRA). Fully fine-tuning LLMs (cf. Equation (2)) entails substantial computational resources [6, 7, 8, 9, 10, 11, 12]. LoRA emerges as an efficient alternative [37, 38, 39, 40, 41, 42, 43], which injects trainable low-rank matrices into transformer layers to approximate the updates of pre-trained weights [42]. At the core, LoRA employs a low-rank decomposition where the update $\Delta\mathbf{W}$ to the pre-trained matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ is represented as $\Delta\mathbf{W}=\mathbf{BA}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and are tunrable up- and down-projection matrices, respectively. The rank $r$ is significantly smaller than both $d_{\text{in}}$ and $d_{\text{out}}$, enhancing adaptation efficiency.
@@ 4 Experiments
In this section, we first justify the need to reshape the fine-tuning task with a uniform LoRA module as a multi-task learning framework for sequential recommendation. Here we conduct extensive experiments on various real-world datasets, including LastFM [31], MovieLens [32], and Steam [33], to evaluate the effectiveness of our iLoRA framework. Our analysis includes detailed comparisons of iLoRA against established baseline models, which encompass both traditional sequential recommender models (e.g., GRU4Rec [1], Caser [2], SASRec [3]) and LLM-based recommender models (e.g., Llama2-7B [9], GPT-4 [44], MoRec [34], TALLRec [16], LLaRA [21]). ValidRatio [21] and HitRatio@1 are used as evaluation metrics, to separately quantify the ratios of valid responses over all sequences and relevant items over all candidate items, reflecting the model capability of instruction following and recommendation accuracy. See Appendix A for more details of these baselines, datasets, and metrics. Moreover, we perform a thorough ablation study to identify the key components that enhance iLoRA’s performance, focusing particularly on the role of the gating network and expert settings. In a nutshell, we would like to answer the following research questions:
@@ Appendix B Experimental Design and Evaluation
Datasets. To validate the generalization ability of iLoRA, we conducted extensive experiments on four datasets derived from real-world recommendation scenarios: LastFM, GoodReads, Amazon-book, and Steam:
  • LastFM Collected from the Last.fm online music platform, including user-artist listening relationships and the names of artists.
  • MovieLens A commonly used movie recommendation dataset that contains user ratings and movie titles.
  • Amazon-book Contains user reviews and interactions with books available on the Amazon platform.
  • Steam Encompasses user reviews for video games on the Steam Store, in addition to game titles.
Baselines. We compared iLoRA with several models, including traditional sequential recommendation models and those based on Large Language Models (LLMs), such as GRU4Rec[1], Caser, SASRec[3], Llama2[9], GPT-4, MoRec, TallRec[16], and LlaRA[21]. GRU4Rec, Caser, and SASRec utilize recurrent neural networks, convolutional neural networks, and Transformer encoders, respectively, to capture sequential patterns in user behavior. Llama2, a well-known open-source LLM, is released by Meta, and in our experiments, we utilized the 7B version. GPT-4, the latest version of the GPT series, is a milestone in the performance of LLMs across diverse tasks, from natural language processing to other domains. We conducted our research using the API provided by OpenAI. MoRec enhances traditional sequential recommendation models, such as text-based features, by leveraging pretrained modal encoders to strengthen the modal characteristics of encoded items. TallRec guides LLMs through fine-tuning recommendation corpora. LlaRA further improves the recommendation effectiveness of LLMs by aligning collaborative signals to the text space through a hybrid prompt method.
Training Protocol. For our study, we opt for Llama2-7B [9] as the LLM backbone. Given the disparity in computational resources between traditional sequential recommendation models and LLM-based methods, we execute baseline models on a single Nvidia A40, while LLaRA is implemented on a single Nvidia-smi a100. This ensures optimal utilization of available computational resources. All implementations are carried out using Python 3.8 and PytorchLightning 1.8.6.
@@ Appendix D Related Work
Large Language Models Recent years have witnessed a surge of activity in language modeling research, establishing it as a cornerstone for both understanding and generating language. This momentum has given rise to a new breed of language models (LMs), including notable works such as BERT [6], GPT-3 [7], LLama [8], LLama2 [9], Mistral-7B [10], Alpaca [11], and Vicuna [12]. These LMs, predominantly based on the Transformer architecture, have demonstrated remarkable versatility, exemplified by models like BERT [6] and T5 [36], owing to their extensive training corpus. A significant stride in this domain has been the exploration of scaling effects, with researchers pushing the boundaries by augmenting both the parameter and training corpus scales to unprecedented magnitudes, encompassing billions of parameters and trillions of training tokens [8, 9, 12, 7, 46]. These Large Language Models (LLMs) exhibit substantial performance enhancements and showcase unique capabilities, including but not limited to common sense reasoning and instruction following. Moreover, the development of domain-specific LLMs further enriches this landscape. Models tailored to specific domains, such as finance [47], medicine [48], and law [49], amalgamate domain expertise with the inherent commonsense knowledge of general LLMs. These advancements not only broaden the scope of LLM applications but also inspire exploration into their potential utility in recommendation systems.
@@ Appendix E Statistics
For all conventional sequential recommendation baselines, we employ the Adam optimization algorithm, establishing a learning rate of 0.001, an embedding dimension d of 64, and a batch size of 256. Furthermore, we implement L2 regularization, with the coefficient determined through a grid search within the range [1e-3, 1e-4, 1e-5, 1e-6, 1e-7]. To mitigate the impact of randomness, we report the mean of the outcomes derived from random seeds within the set [0, 1, 2, 3, 4]. For all methods related to LLMs, we employ a warm-up strategy for the learning rate. This strategy initiates the training process with a relatively modest learning rate, set at 1/100 of the maximum learning rate. The max learning rates are 2e-4 and 1e-4 for the LastFM, MovieLens and Steam datasets, respectively. Note that the parameters of the projector are updated during the training process using the same learning rate curve with LoRA. Additionally, we utilize a cosine scheduler to adjust the learning rate over steps, and we implement half-precision computation to optimize memory usage and computational efficiency during the experiments. Each experiment is trained for a maximum of five epochs, with a batch size of 128. Furthermore, we apply a weight decay factor of 1e-5 to prevent overfitting.

### v3

@@ Abstract
Sequential recommendation systems predict the next interaction item based on users’ past interactions, aligning recommendations with individual preferences. Leveraging the strengths of Large Language Models (LLMs) in knowledge comprehension and reasoning, recent approaches are eager to apply LLMs to sequential recommendation. A common paradigm is converting user behavior sequences into instruction data, and fine-tuning the LLM with parameter-efficient fine-tuning (PEFT) methods like Low-Rank Adaption (LoRA). However, the uniform application of LoRA across diverse user behaviors is insufficient to capture individual variability, resulting in negative transfer between disparate sequences. To address these challenges, we propose Instance-wise LoRA (iLoRA). We innovatively treat the sequential recommendation task as a form of multi-task learning, integrating LoRA with the Mixture of Experts (MoE) framework. This approach encourages different experts to capture various aspects of user behavior. Additionally, we introduce a sequence representation guided gate function that generates customized expert participation weights for each user sequence, which allows dynamic parameter adjustment for instance-wise recommendations. In sequential recommendation, iLoRA achieves an average relative improvement of 11.4% over basic LoRA in the hit ratio metric, with less than a 1% relative increase in trainable parameters. Extensive experiments on three benchmark datasets demonstrate the effectiveness of iLoRA, highlighting its superior performance compared to existing methods in mitigating negative transfer and improving recommendation accuracy. Our data and code are available at https://github.com/AkaliKong/iLoRA.
@@ 1 Introduction
Sequential recommendation [1] suggests a user’s next item of interest by analyzing his/her past interactions, tailoring recommendations to individual preferences. As Large Language Models (LLMs) [2] exhibit impressive proficiency in global knowledge comprehension and reasoning, their potential for application in sequential recommendation is garnering increasing interest [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Recent efforts [9, 10] approach the sequential recommendation task under a language generation paradigm, wherein user behavior sequences are converted into input prompts by either purely textual prompting (ID numbers or descriptions) [3, 5, 15, 6] or hybrid prompting with additional behavioral tokens [9, 10, 16], achieving remarkable success. Upon scrutinizing prior studies on LLM-based sequential recommenders, we can summarize a common fine-tuning pipeline comprising three components: (1) convert a sequence of historical behaviors into a prompt; (2) pair these prompts with the subsequent items of interest, to create instruction-tuning datasets; (3) incorporate a trainable Low-Rank Adaptation (LoRA) module [5, 16, 10, 9] into LLMs and fine-tune it on such prompts. Existing studies [17, 3, 18, 6, 9, 10, 16] primarily focus on refining high-quality prompts to more effectively incorporate recommendation information, while leaving the choice of LoRA unexplored. Instead, they employ a standard LoRA module for fine-tuning, which freezes the LLM weights and updates the model through two additional low-rank matrices.
@@
Previous work [19, 20] have demonstrated that related tasks tend to develop similar loss geometries, while unrelated tasks exhibit dissimilar ones. Using the same set of parameters in a multi-task learning context can lead to conflicts, especially when tasks have low gradient similarity. This can result in negative transfer, which limits further improvement of the model. From this perspective, since user behaviors often exhibit substantial individual variability (e.g., distinct interests, behavior patterns, feedback mechanisms), we argue that employing a standard LoRA module across such diverse behaviors may not effectively capture these variabilities. Specifically, the inherent variability in user behaviors naturally inspires us to view item sequences with different behavior variables as different tasks. Simply applying a singular LoRA module leaves this multitask nature of sequential recommendation untouched, thus potentially overestimating the relationships between sequences. It easily causes the negative transfer between significantly discrepant sequences. Take LLaRA [9] as an example, which fine-tunes a LoRA module on top of Llama-2 [21] across all sequences. Following prior studies [19, 20], we use a symmetric heatmap to analyze the pair-wise gradients similarities of LLaRA associated with disparate sequences and reveal significant misalignment, as shown in Figure 1. Specifically, we observe strong clustering by membership closeness in the collaborative space, along the diagonal of the gradient similarity matrix. Meanwhile, clusters that are more distant in the collaborative space tend to exhibit dissimilar gradient trajectories. Clearly, such gradients are misaligned, which can result in suboptimal performance. To mitigate this issue, one straightforward solution is to deploy multiple LoRA modules, each fine-tuned for a specific sequence, enabling each module to act as a lightweight expert tailored to its respective sequence. However, it is impractical in terms of resources and time complexity, as the number of sequences often scales to millions.
@@
Figure 1: Gradient similarity of LoRA modules across training steps. The sequence dataset is partitioned into 8 clusters using Euclidean distance, with hierarchical clustering applied to reorder clusters, so that clusters closer in the collaborative space are also closer together in the heatmap. Gradient similarity is used to assess the geometric characteristics of the loss, with darker cells indicating higher similarity. In the case study on the right, dashed lines connect similar items, while solid lines link identical items. Users with a gradient similarity of 0.86 share a strong interest in thriller movies, while those with -0.75 cosine similarity show no clear preference alignment.
@@ 2 Preliminary
Fine-tuning with Low-rank Adaption (LoRA). Fully fine-tuning LLMs (cf. Equation (2)) entails substantial computational resources [36, 35, 37, 21, 38, 39, 40]. LoRA emerges as an efficient alternative [41, 42, 43, 44, 45, 46, 47, 48, 49], which injects trainable low-rank matrices into transformer layers to approximate the updates of pre-trained weights [46]. At the core, LoRA employs a low-rank decomposition where the update $\Delta\mathbf{W}$ to the pre-trained matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ is represented as $\Delta\mathbf{W}=\mathbf{BA}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and are tunrable up- and down-projection matrices, respectively. The rank $r$ is significantly smaller than both $d_{\text{in}}$ and $d_{\text{out}}$, enhancing adaptation efficiency.
@@ 4 Experiments
In this section, we first justify the need to reshape the fine-tuning task with a uniform LoRA module as a multi-task learning framework for sequential recommendation. We request the dataset from LLaRA [9] and maintain exactly the same experimental settings as described in the original paper. Our study builds upon the main table from the LLaRA paper, using the results reported in the LLaRA paper directly. Here we conduct extensive experiments on various real-world datasets, including LastFM [25], MovieLens [26], and Steam [27], to evaluate the effectiveness of our iLoRA framework. Our analysis includes detailed comparisons of iLoRA against established baseline models, which encompass both traditional sequential recommender models (e.g., GRU4Rec [28], Caser [29], SASRec [24]) and LLM-based recommender models (e.g., Llama2-7B [21], GPT-4 [50], MoRec [30], TALLRec [5], LLaRA [9]). ValidRatio [9] and HitRatio@1 are used as evaluation metrics, to separately quantify the ratios of valid responses over all sequences and relevant items over all candidate items, reflecting the model capability of instruction following and recommendation accuracy. See Appendix A for more details of these baselines, datasets, and metrics. Moreover, we perform a thorough ablation study to identify the key components that enhance iLoRA’s performance, focusing particularly on the role of the gating network and expert settings. In a nutshell, we would like to answer the following research questions:
@@ 7 Broader Impact
Acknowledgments and Disclosure of Funding
This research is supported by the National Natural Science Foundation of China (92270114, 62302321) and the National Science and Technology Major Project (2023ZD0121102). This research is also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Appendix B Experimental Design and Evaluation
Datasets. To validate the generalization ability of iLoRA, we conducted extensive experiments on three datasets derived from real-world recommendation scenarios: LastFM, MovieLens, and Steam:
  • LastFM A popular music dataset for recommendation research, featuring histories of user-artist interactions, user demographic details, artists’ names and associated social tags.
  • MovieLens A widely used benchmark in recommendation systems, containing user ratings and metadata for movies.
  • Steam A collection of user interaction data from the Steam gaming platform, featuring information on game ownership, playtime, and user reviews.
Baselines.
We compared iLoRA with several models, including traditional sequential recommendation models and those based on Large Language Models (LLMs), such as GRU4Rec[28], Caser, SASRec[24], Llama2[21], GPT-4, MoRec, TallRec[5], and LlaRA[9]. GRU4Rec, Caser and SASRec utilize recurrent neural networks, convolutional neural networks, and Transformer encoders, respectively, to capture sequential patterns in user behavior. Llama2, Meta’s open-source language model, building on the original Llama to deliver improved performance. We fine-tune the 7B version of the model for our experiments. GPT-4, OpenAI’s advanced language model, has held the state-of-the-art position across various tasks for an extended period. We directly utilize its API for our experiments. MoRec uses pre-trained modality encoders to capture item-specific features, improving recommendation accuracy. TALLRec guides LLMs through fine-tuning recommendation corpora. LlaRA further improves the recommendation effectiveness of LLMs by aligning collaborative signals to the text space through a hybrid prompt method.
Training Protocol.
In our study, we fine-tune the Llama2-7B [21] model to validate our approach. Experiments for traditional sequential recommendation baseline models are conducted on a single Nvidia A40, while our iLoRA framework is implemented on a single Nvidia A100. All experiments are carried out using Python 3.8 and PytorchLightning 1.8.6.
@@ Appendix D Related Work
Large Language Models Sequential recommendation [24, 32, 31, 52, 53, 54] aims to predict the user’s next likely item of interest by analyzing their past interaction patterns and aligning with their preferences. Recent years have witnessed a surge of activity in language modeling research, establishing it as a cornerstone for both understanding and generating language. This momentum has given rise to a new breed of language models (LMs), including notable works such as BERT [36], GPT-3 [35], LLama [37], LLama2 [21], Mistral-7B [38], Alpaca [39], and Vicuna [40]. These LMs, predominantly based on the Transformer architecture, have demonstrated remarkable versatility, exemplified by models like BERT [36] and T5 [34], owing to their extensive training corpus. A significant stride in this domain has been the exploration of scaling effects, with researchers pushing the boundaries by augmenting both the parameter and training corpus scales to unprecedented magnitudes, encompassing billions of parameters and trillions of training tokens [37, 21, 40, 35, 55]. These Large Language Models (LLMs) exhibit substantial performance enhancements and showcase unique capabilities, including but not limited to common sense reasoning and instruction following. Moreover, the development of domain-specific LLMs further enriches this landscape. Models tailored to specific domains, such as finance [56], medicine [57], and law [58], amalgamate domain expertise with the inherent commonsense knowledge of general LLMs. These advancements not only broaden the scope of LLM applications but also inspire exploration into their potential utility in recommendation systems.
@@ Appendix E Statistics
In our experimental settings, we closely follow the setup of prior works [9] to ensure a fair and meaningful comparison. Specifically, for all conventional sequential recommendation baselines, we employ the Adam optimization algorithm, establishing a learning rate of 0.001, an embedding dimension of 64, and a batch size of 256, respectively. Furthermore, we incorporate L2 regularization, and the regularization coefficient is fine-tuned through a grid search over a specified set of possible values.We report the average results from five different random seeds, specifically selected from the set, To reduce the effect of randomness.
For experiments involving methods based on large language models (LLMs), we incorporate a warm-up strategy for the learning rate, which begins the training process with an initial rate set at a fraction of the maximum learning rate. Specifically, the maximum learning rates are set to $2e^{-4}$ for the LastFM dataset and $1e^{-4}$ for the MovieLens and Steam datasets. Notably, the projector parameters are updated using the same learning rate curve integrated with iLoRA during training.
We adopt a cosine annealing scheduler to adjust the learning rate across training steps, which facilitates a smooth decrease in the learning rate, thereby improving stability during training. Additionally, half-precision computation (mixed-precision training) is used to enhance memory efficiency and accelerate the training process.